Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Ring Heater)
nardecchia, menzione - 1:22 Saturday 24 February 2024 (63391) Print this report
NE RH WP investigation

On Sunday 18th Feb, Andrea settled the RH powers as in O3b (63309):

  O3b values

NE

WE

18th Feb

(14.19 UTC)

4.7 W

[V=13.2 V]

6.3 W 

[V=15.3 V]

For the  NE RH, the O3b power (5.1 W)  has been re-scaled considering that the cold RoC of the new NE mirror is 1692.7 m, ~3.3 m lower than the old one (1696 m).

The afternoon shift has been dedicated to investigate the possibility to improve the  NE RH  WP.

LN3 has been reached at  15.48 UTC.

Diego and Manuel completed their shift around 17 UTC. 

The NE RH power has been increased by 0.2 W at 17.02 UTC:

4.7 W (V=13.2 V) → 4.9 W (V=13.5 V)

19.13 UTC: unlock

The main ITF signals during this lock are shown in fig.1 

It was impossible to relock due to different problems reported in 63389.

Nicola solved them but, unfortunately the ITF unlocked systematically in the first few minutes after reaching carm null 1f without particular oscillations.

To investigate the issue, the NE RH working point has been changed decreasing the power from 4.9 W to 4.8 W (V=13.32 V) at 22.22 UTC.

Then, I waited some time in CITF profiting to adjust the CH powers to optimize the sidebands levels. The new values are in bold in the table below. 

 

REFERENCE

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.039 -->0.047

0

1.27

on the pickoff 

0.242 -->0.291

0

0.208

N

on the ITF

0.106 -->0.116

0.34

3.68

on the pickoff 

0.655 -->0.715

0.055

0.60

Then, the lock acquisition re-started and, finally,  at 23.38 UTC, the ITF reached LN2 and at 23.50 UTC LN3 (but with a bad sensitivity).

The main signals during the whole shift are shown in fig.2.

Images attached to this report
Search Help
×

Warning

×